OpenStudy (anonymous):

There are 70 teachers in a school. Of the 38 male teachers, 8 teach Maths It is known that 2/7 of all teachers teach Maths. Find the probability that a teacher selected at random is. (a) a male teacher or a teacher who teaches Maths. (b) a female teacher or a teacher who teaches Maths. (c) a female teacher or a teacher who does not teach Maths. @ganeshie8

OpenStudy (anonymous):

2/7 of 70 gives 20 that means ...20 maths teachers ..

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

8 are male maths teacher and 12 are female maths teacher

OpenStudy (anonymous):

oh yes.

OpenStudy (anonymous):

then for part a ..what will u do ? p(E)=favourable /total outcomes

OpenStudy (anonymous):

8/38 + 20/70 ???

OpenStudy (anonymous):

no ...first is a male teacher ...38/70 and then a teacher who teaches maths will be 2/7 ...sum of these two probabilitites by 2

OpenStudy (anonymous):

by 2 because at random also we should get either a male teacher or a teacher who teaches maths

OpenStudy (anonymous):

addition isn't it?

OpenStudy (anonymous):

\[\frac{ 38/70+2/7 }{ 2 }\]

OpenStudy (anonymous):

for part b\[\frac{ 32/70+2/7 }{ 2}\]

OpenStudy (anonymous):

for part c\[\frac{ 32/70+4/7 }{ 2 }\]

OpenStudy (anonymous):

wait, i don't think part a is correct. the answer is 5/7

OpenStudy (anonymous):

ohhhh ....yes sorry i went somewhere .wrong in understanding ...!

OpenStudy (anonymous):

this is addition law of probability, isn't it?

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

does the question : "a male teacher or a teacher who teaches Maths" means "a male teacher who is not exactly teaches Maths or a teacher who teaches Maths" ?

OpenStudy (anonymous):

apply and see if the answer comes 5/7 ..? den interpretation may be right .

OpenStudy (anonymous):

interpretation ?